40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1374     Entry time: Mon Mar 9 12:04:18 2009
Author: Yoichi 
Type: Update 
Category: Computers 
Subject: TPs and AWG are back 
I had to do one more reboot of tpman and daqd to get the TPs working.
I confirmed the alignment scripts run fine.

Now the oplevs of some optics are largely mis-centered. Alberto and I will center them after lunch.
ELOG V3.1.3-